# Beyond Reference-Based Metrics: Analyzing Behaviors of Open LLMs on Data-to-Text Generation

### v1

@@ Abstract
We investigate to which extent open large language models (LLMs) can generate coherent and relevant text from structured data. To prevent bias from benchmarks leaked into LLM training data, we collect Quintd-1: an ad-hoc benchmark for five data-to-text (D2T) generation tasks, consisting of structured data records in standard formats gathered from public APIs. We leverage reference-free evaluation metrics and LLMs’ in-context learning capabilities, allowing us to test the models with no human-written references. Our evaluation focuses on annotating semantic accuracy errors on token-level, combining human annotators and a metric based on GPT-4. Our systematic examination of the models’ behavior across domains and tasks suggests that state-of-the-art open LLMs with 7B parameters can generate fluent and coherent text from various standard data formats in zero-shot settings. However, we also show that semantic accuracy of the outputs remains a major issue: on our benchmark, 80% of outputs of open LLMs contain a semantic error according to human annotators (91% according to GPT-4). Our code, data, and model outputs are available at https://d2t-llm.github.io.
@@ 1 Introduction
Figure 1: We experiment with using LLMs for generating text from structured data in various domains, analyzing model behavior and evaluating their output on token level.
@@
wikidata World facts Desribing entities and relations in a knowledge graph. Wikidata Markdown
@@
Table 1: The domains and tasks included in the Quintd data collection tool we use for testing D2T generation with LLM. In our experiments, we download 100 development and 100 test examples of input data for each domain.
@@
Large language models (LLMs; Ouyang et al., 2022; Touvron et al., 2023a, b; Jiang et al., 2023; Tunstall et al., 2023) have already left a mark in many areas of natural language processing (NLP). Suprisingly, their applicability to the task of data-to-text (D2T) generation Reiter and Dale (1997); Gatt and Krahmer (2018) remains underexplored, with limited evaluation on a handful of well-established benchmarks only Axelsson and Skantze (2023); Yuan and Färber (2023). Generating text from structured data is arguably challenging for LLMs, given the specifics of D2T generation such as long inputs, complex non-linear structure, and strict requirements on semantic accuracy. However, a bigger issue is perhaps the lack of testing grounds. The current D2T generation benchmarks are saturated by approaches optimizing towards traditional reference-based evaluation metrics, which were shown to correlate poorly with human judgement Gehrmann et al. (2023); van der Lee et al. (2021); Novikova et al. (2017). When it comes to the models, using closed LLMs OpenAI (2023a, b) is increasingly considered a bad research practice due to its non-reproducibility Rogers (2023); Chen et al. (2023), and potential data contamination further restricts the space for experiments Golchin and Surdeanu (2023); Aiyappa et al. (2023); Anonymous (2023).
@@
In this paper, we propose how to get around these issues and move towards better D2T generation approaches with more rigorous evalution. We begin with the realization that unlabeled data are plentiful. To leverage the data for our experiments, we introduce Quintd: [^1] Quintet of Unlabeled Inputs for Natural Tasks in Data-to-text (pronounced as “quintet”). a tool for collecting structured data from five domains in standard formats: JSON, CSV, and Markdown. We choose the domains so that the data can be directly used as input for five distinct D2T generation tasks. Specifically, our tasks inlude generating weather forecasts, sports reports, product descriptions, chart captions, and entity descriptions (see Table 1). Next, we collect a set of 1000 inputs with Quintd and use the inputs as an ad-hoc benchmark (called Quintd-1) for testing abilities of LLMs for D2T generation. We assume that the data formats in Quintd-1 are common in the LLMs’ pretraining corpora, so we specify the task using instructions (instead of standard finetuning with human-written outputs), capitalizing on the in-context learning abilities of LLMs (§2).
@@
We push towards better reproducibility by focusing on open LLMs, which – apart from being more accessible – also achieve increasingly better results across tasks Zheng et al. (2023); Beeching et al. (2023). For our experiments, we use three open LLMs with 7B parameters: Llama-2 Touvron et al. (2023b); TogetherAI (2023), Mistral Jiang et al. (2023), and Zephyr Tunstall et al. (2023). We also use GPT-3.5 OpenAI (2023b) as a closed model baseline for the final experiments. Given the behavioral nature of the experiments with LLMs Holtzman et al. (2023), we put emphasis on reporting model behavior throughout the process (§3).
@@
Another piece of the puzzle is reference-free evaluation: using the input data as a ground for comparison, instead of human references (§4). For evaluation, we use manual annotations from human crowdworkers van der Lee et al. (2021), along with a customized automatic metric based on GPT-4 Liu et al. (2023); Chiang and Lee (2023); Kocmi and Federmann (2023a). To get a fine-grained picture of model errors, we annotate semantic accuracy errors on the level of individual tokens Thomson and Reiter (2020); Thomson et al. (2023).
@@
Based on our results, we provide general recommendations about D2T generation with open LLMs across tasks and formats (§5). Our main findings are as follows:
@@
  • Semantic accuracy is still a major obstacle: on average, 80% of LLM outputs contain a semantic error according to human annotators and 91% according to the GPT-4-based metric.
@@
  • The output quality can be aided by following several rules-of-thumb on preprocessing the model input, such as including units, removing unnecessary fields, or prefixing the model answer.
@@ 2.1 Data Collection Tool
We introduce a tool dubbed Quintd for collecting ad-hoc tests sets using public APIs in five different domains, and we collect one such set (Quintd-1) for our experiments. Our main reasons for departing from the traditional scheme of benchmarking on well-established datasets are:
@@
Start with
Figure 2: The prompt $\mathcal{P}$ and the model output prefix (“Start with”) we used for the experiments in this paper. DATA is filled with the data record $x$ and output_type is filled accordingly for each domain $\mathcal{D}$ (cf. Table 1).
@@ 2.2 Quintd-1 Dataset
The benchmark we collected using Quintd for our experiments in this paper (Quintd-1) contains 500 examples in the development set and 500 examples in the test set (100 examples per domain for each split). We keep the size of the dataset moderate for a quick experimental turnaround.
We downloaded the data between November 2023 and January 2024. Note that the dataset contains only unlabeled data without any reference outputs (e.g. weather data, but not a textual weather forecast). New versions of the benchmark can be easily created with the Quintd tool we provide.
@@ 3.1 Experimental Process
Our goal is to avoid extensive data preprocessing and prompt engineering, since these steps could harm the reproducibility and generalizability of our experiments. With this goal in mind, we decided to use the same prompt template $\mathcal{P}$ for all the domains and models.
For a set of preliminary experiments, we first wrote down the initial version of the prompt and used the data without further preprocessing. We then iteratively improved our experimental setup by observing outputs on the development set. We describe all the observations and modifications we made before generating the final outputs on the test set in §3.3.
Figure 3: Our experimental setup. We first generate the outputs using LLMs which are given raw data and a task-specific prompt. We annotate the token-level semantic errors in the LLM outputs with a) an automatic metric based GPT-4 which matches the output to the raw data, and b) human annotators, who annotate the errors in the output given the data visualization.
@@ 3.2 Models
The models are instruction-tuned, operate with 32k context, and have good performance on recent benchmarks. All the models have 7B parameters and thus fit on a single NVIDIA A40 (48G VRAM) in 16-bit precision. The models are available through HuggingFace Wolf et al. (2020).
We accessed the models via an API provided by the text-generation-webui framework [^2] https://github.com/oobabooga/text-generation-webui running locally. For the final experiments, we also included GPT-3.5 OpenAI (2023b) accessed through the OpenAI API.
@@ 3.3 Observations from Preliminary Experiments
The models do not always select the most relevant fields for the given output. For example, we observed that the models commonly mention identifiers, timestamps, files, and other metadata, leading to unnatural outputs. Due to this, we decided not to include these irrelevant fields on the input.
@@
If the units are not specified in the data record, the models tend to resort to their best guess. This may go unnoticed if the unit is clear from the context (e.g., the model will usually not report temperature in Fahrenheit instead of Celsius), but it may get problematic if the value is ambiguous (e.g., wind speed in km/h versus m/s). Therefore, we explicitly add units to all data records where appropriate.
Understandable labels are enough.
On the flip side, we decide not to add extra descriptions to the keys if the key was understandable from the label (e.g. homeTeam or dimensions). As discussed by Kasner et al. (2023b), pretrained models tend to interpret the fields correctly as long as the label is human-readable. We only decided to include chart metadata for the CSV files in the owid domain.
@@
The inputs in some domains can easily get longer than 10-20k tokens. This issue is amplified by the fact that the evaluated LLMs tokenize numbers into individual digits. Because of the long inputs, we decided to give up the official Llama-2-7B-Instruct model with 4k context Touvron et al. (2023b) in favor of Llama2 with 32k context TogetherAI (2023). With long inputs, the GPU memory consumption also gets considerably higher, so we needed to downsample the data in owid and openweather to keep their length under ~8k tokens.
@@
Due to the context-length limitations described above, we were not able to conduct few-shot experiments since we could not robustly fit an additional ($x_{\text{example}}$, $y_{\text{example}}$) pair in the prompt. We attempted including only $y_{\text{example}}$ (making the setup “half-shot”), but we observed that the models then tended to use entities from the example (unrelated to the actual input) in their outputs. Therefore, we decided not to to follow this line of experiments (see §5.3 for discussion).
@@
In our preliminary experiments, we observed a roughly similar output quality for both deterministic decoding and sampling. [^3] We used the text-generation-webui default decoding parameters: temperature=0.7, top_p=0.9, and top_k=20. For the final experiments, we decided to use deterministic decoding which is non-parametric and conceptually more suitable for D2T generation.
@@
The outputs are fluent, but inaccurate.
We observed that the vast majority of model outputs were gramatically and stylistically correct, capturing the output type specified in the prompt. However, we also noticed that the outputs contained many factual errors (even after emphasizing the focus on factual accuracy in the prompt, see Figure 2). This observation lead us to evaluate the model outputs using token-level annotations focused on semantic accuracy errors Reiter and Thomson (2020).
@@ 3.4 Final Experiments
Taking the observations in §3.3 into account, we proceeded to generate the outputs on the test set of Quintd-1 for token-level error analysis. We first preprocessed the data as mentioned: we stripped out unnecessary fields, added units, and downsampled the data to fit the context. For all the models mentioned in §3.2, we used the prompt in Figure 2 and deterministic decoding with a maximum length of 512 tokens.
For comparsion, we also generated outputs for the same inputs and identical prompts with GPT-3.5 (gpt-3.5-turbo-1106). [^4] We did not include GPT-3.5 in our preliminary experiments since closed models are not our focus. We also did not use GPT-4 because we reserve the model for evaluation, see §4.1. Given that the decoding parameters provided by the OpenAI API are fairly limited, we only tried to bring the setup closer to deterministic decoding by fixing the temperature and seed to 00.
@@ 4 Evaluation
For evaluation, we focus on semantic accuracy errors. We compare the generated texts to the input data, looking for parts of texts which are not faithful to the input data. We annotate the errors on the token level, considering all the tokens in the output text as potential sources of errors.
Regarding the error taxonomy, we settled on four error categories: INCORRECT, NOT_CHECKABLE, MISLEADING, and OTHER. The taxonomy is inspired by the methodology discussed in Thomson and Reiter (2020) and Thomson et al. (2023). To keep the annotation tractable, we decided not to distinguish between fine-grained categories (e.g. incorrect name vs. incorrect number). The descriptions of our error categories, as presented in the instructions for annotation, are included in Table 2.
@@
INCORRECT
The fact in the text contradicts the data.
NOT_CHECKABLE
The fact in the text cannot be checked given the data.
MISLEADING
The fact in the text is misleading in the given context.
OTHER
The text is problematic for another reason, e.g. grammatically or stylistically incorrect, irrelevant, or repetitive.
@@
data
Nokia 3310 | color: black, blue, grey | display: 320x240px
text
Nokia 3310 is produced in Finland and features a 320x320 display. It is available in black color. The data seem to provide only partial information about the phone.
Table 2: Categories of errors annotated in our evaluation and an example demonstrating the error types. See Appendix C for explanation about individual errors in the example.
We employ two complementary evaluation schemes:
  • $\mathcal{E}_{\text{gpt}}$: an automatic metric based on GPT-4 (§4.1),
  • $\mathcal{E}_{\text{hum}}$: human evaluation based on crowdsourcing (§4.2).
These two schemes are based on similar instructions and produce (nearly) equivalent outputs. The main idea in introducing multiple schemes is to compensate for shortcomings of each approach and thus increase the replicability and robustness of our results.

### v3

# Beyond Traditional Benchmarks: Analyzing Behaviors of Open LLMs on Data-to-Text Generation

@@ Abstract
We analyze the behaviors of open large language models (LLMs) on the task of data-to-text (D2T) generation, i.e., generating coherent and relevant text from structured data. To avoid the issue of LLM training data contamination with standard benchmarks, we design Quintd – a tool for collecting novel structured data records from public APIs. We find that open LLMs (Llama 2, Mistral, and Zephyr) can generate fluent and coherent texts in zero-shot settings from data in common formats collected with Quintd. However, we show that the semantic accuracy of the outputs is a major issue: both according to human annotators and our reference-free metric based on GPT-4, more than 80% of the outputs of open LLMs contain at least one semantic error. We publicly release the code, data, and model outputs. [^1] https://d2t-llm.github.io/
@@ 1 Introduction
Figure 1: To benchmark LLMs, we download unlabeled structured data from public APIs and prompt LLMs to generate texts based on the data. We annotate semantic errors in the outputs using reference-free metrics.
@@
wikidata World facts Describing entities and relations in a knowledge graph. Wikidata Markdown
@@
Table 1: The domains and tasks included in the Quintd data collection tool we use for testing D2T generation with LLMs. In our experiments, we download 100 development and 100 test examples of input data for each domain.
@@
Large language models (LLMs; Ouyang et al., 2022; Touvron et al., 2023a, b; Jiang et al., 2023; Tunstall et al., 2023) have already left a mark in many areas of natural language processing (NLP). Surprisingly, their applicability to the task of data-to-text (D2T) generation Reiter and Dale (1997); Gatt and Krahmer (2018) remains underexplored, with limited evaluation on a handful of well-established benchmarks only Axelsson and Skantze (2023); Yuan and Färber (2023). Generating text from structured data is arguably challenging for LLMs, given the specifics of D2T generation, such as long inputs, complex non-linear structure, and strict requirements on semantic accuracy. However, a more significant issue is the lack of testing grounds. The current D2T generation benchmarks are not only getting saturated Van Miltenburg et al. (2023), but also promote optimization towards traditional reference-based evaluation metrics, which were shown to correlate poorly with human judgment Gehrmann et al. (2023); van der Lee et al. (2021); Novikova et al. (2017). When it comes to the models, using closed LLMs OpenAI (2023a, b) is increasingly considered a bad research practice due to its non-reproducibility Rogers (2023); Chen et al. (2023). On top of that, contamination of LLM training data with standard benchmarks further restricts the space for experiments Golchin and Surdeanu (2023); Aiyappa et al. (2023); Balloccu et al. (2024).
@@
In this paper, we propose an approach that allows us to analyze model behavior in D2T generation on novel, real-world structured data records with reference-free evaluation metrics. We begin by realizing that unlabeled data are plentiful. To leverage the data for our experiments, we introduce Quintd  [^2] Quintet of Unlabeled Inputs for Natural Tasks in Data-to-text, pronounced as “quintet” – a tool for collecting structured data from five domains in standard formats: JSON, CSV, and Markdown. We choose the domains so that the data can be directly used as input for five distinct D2T generation tasks. Our tasks include generating weather forecasts, sports reports, product descriptions, chart captions, and entity descriptions (see Table 1). Next, we collect a set of 1,000 inputs with Quintd and use the inputs as an ad-hoc benchmark (called Quintd-1) for testing the abilities of LLMs for D2T generation. We assume that the data formats in Quintd-1 are common in the LLMs’ pretraining corpora, so we specify the task using instructions instead of standard finetuning with human-written outputs, capitalizing on the zero-shot abilities of instruction-tuned LLMs (§2).
@@
We push towards better reproducibility by focusing on open LLMs, which – apart from being more accessible – also achieve increasingly better results across tasks Zheng et al. (2023); Beeching et al. (2023). For our experiments, we use three open LLMs with 7B parameters: Llama 2 Touvron et al. (2023b); TogetherAI (2023), Mistral Jiang et al. (2023), and Zephyr Tunstall et al. (2023). We also use GPT-3.5 OpenAI (2023b) as a closed model baseline for the final experiments. Given the behavioral nature of the experiments with LLMs Holtzman et al. (2023), we put emphasis on reporting model behavior throughout the process (§3).
@@
Another piece of the puzzle is reference-free evaluation: using the input data as a ground for comparison instead of reference outputs (§4). We focus on identifying semantic errors in the model outputs, i.e., the information that is not supported by the input data. We use two separate evaluation methods: manual annotations from human crowdworkers van der Lee et al. (2021) and a custom automatic metric based on GPT-4 Liu et al. (2023); Chiang and Lee (2023); Kocmi and Federmann (2023a). We annotate the errors on the level of individual words, getting fine-grained annotations of error spans in several categories Thomson and Reiter (2020); Thomson et al. (2023).
@@
Based on our results, we provide general recommendations for D2T generation with open LLMs across tasks and formats (§5). Our main findings are as follows:
@@
  • Semantic accuracy is a major obstacle: both human annotators and GPT-4-based metric report that over 80% of outputs of open LLMs on our data contain a semantic error.
@@
  • Outputs can be empirically improved by following several rules-of-thumb for preprocessing the model input, such as including units, removing unnecessary fields, or prefixing the model answer.
@@ 2.1 Data Collection Tool
We introduce Quintd, [^3] https://github.com/kasnerz/quintd a tool for collecting ad-hoc test sets using public APIs in five different domains. Our main reasons for departing from the traditional scheme of benchmarking on well-established datasets are:
@@
Based on the given data: ``` {data} ``` Your task is to write a brief, fluent, and coherent single-paragraph {output_type} in natural language. The text should be balanced and neutral. Make sure that all the facts mentioned in the text can be derived from the input data, do *not* add any extra information.
Output prefix
Sure! Here is the {output_type}: "
Figure 2: The prompt $\mathcal{P}$ and the model output prefix we used for the experiments in this paper. data is filled with the data record $x$ and output_type is filled accordingly for each domain $\mathcal{D}$ (see Table 1 and Table 6 in the Appendix).
@@ 2.2 Quintd-1 Dataset
Using Quintd, we collected the dataset for our experiments in this paper (Quintd-1). The dataset contains 500 examples in the development set and 500 examples in the test set (100 examples per domain for each split). We downloaded the data between November 2023 and January 2024. Note that the dataset contains only unlabeled data without any reference outputs (e.g., weather data, but not a textual weather forecast), so the outputs need to be evaluated using reference-free metrics. New versions of the benchmark can be easily generated with our Quintd tool.
@@ 3.1 Experimental Process
Our goal is to avoid extensive data preprocessing and prompt engineering since these steps could harm the reproducibility and generalizability of our experiments. With this goal in mind, we decided to use the same prompt template $\mathcal{P}$ for all the domains and models.
For a set of preliminary experiments, we first wrote down the initial version of the prompt and used the data without further preprocessing. We then iteratively improved our experimental setup by observing outputs on the development set. In §3.3, we describe all the observations and modifications we made before generating the final outputs on the test set.
Figure 3: Our experimental setup. We first generate the outputs using LLMs that are given raw data and a task-specific prompt. We annotate the word-level semantic errors in the LLM outputs with (a) an automatic metric based on GPT-4 that matches the output to the raw data, and (b) human annotators, who annotate the errors in the output given the data visualization.
@@ 3.2 Models
The models are instruction-tuned, operate with 32k context, and perform well on recent benchmarks Beeching et al. (2023). All the models have 7B parameters and thus fit on a single NVIDIA A40 (48G VRAM) in 16-bit precision. The models are available through HuggingFace Wolf et al. (2020).
We accessed the models via an API provided by the text-generation-webui framework [^4] https://github.com/oobabooga/text-generation-webui running locally. For the final experiments, we also included GPT-3.5 (gpt-3.5-turbo-1106) accessed through the OpenAI API OpenAI (2023b). [^5] We only included GPT-3.5 in our final experiments as proprietary models were not our focus. We did not include GPT-4 since we use the same model for evaluation (see §4.2) and LLMs tend to be biased towards their own outputs Koo et al. (2023); Stureborg et al. (2024).
@@ 3.3 Observations from Preliminary Experiments
The models do not always select the most relevant fields for the given output. For example, we observed that the models commonly mention identifiers, timestamps, files, and other metadata, leading to unnatural outputs. To mitigate these issues, we manually picked irrelevant fields and filtered them out from the input.
@@
If the units are not specified in the data record, the models tend to resort to their best guess. This may go unnoticed if the unit is evident from the context (e.g., the model will usually not report the temperature in Fahrenheit instead of Celsius), but it may get problematic if the value is ambiguous (e.g., wind speed in km/h versus m/s). Therefore, we explicitly add units to all data records where appropriate.
Understandable field names are enough.
On the flip side, we decided not to add extra descriptions to field names in the data if the field was understandable from its name (e.g., homeTeam or dimensions). As discussed by Kasner et al. (2023), pretrained models interpret field names correctly as long as they are human-readable. We only include chart metadata for the CSV files in the owid domain.
@@
The inputs in some domains can easily get longer than 10-20k tokens. This issue is amplified by the fact that the evaluated LLMs tokenize numbers into individual digits. To accommodate for the long inputs, we picked models that accept up to 32k tokens. [^6] For this reason, we use Llama-2-7B-32k with 32k token context TogetherAI (2023) instead of the official Llama-2-7B-Instruct, which only supports 4k context Touvron et al. (2023b). However, with long inputs, the GPU memory consumption also gets considerably higher, so we needed to downsample the data in owid and openweather to keep their length under ~8k tokens.
@@
Due to the above context-length limitations, we were not able to run few-shot experiments since we could not robustly fit an additional ($x_{\text{example}}$, $y_{\text{example}}$) pair in the prompt. We attempted to include only $y_{\text{example}}$ (making the setup “half-shot”), but we observed that the models then used entities from the example (unrelated to the actual input) in their outputs. Therefore, we decided to leave this line of experiments for future work (see §5.3 for discussion).
@@
In our preliminary experiments, we observed a roughly similar output quality for both deterministic decoding and sampling. [^7] We used the text-generation-webui default decoding parameters: temperature=0.7, top_p=0.9, and top_k=20. For the final experiments, we decided to use deterministic decoding, which is non-parametric and conceptually more suitable for D2T generation.
@@
The outputs are fluent but inaccurate.
We observed that the vast majority of model outputs were grammatically and stylistically correct, capturing the output type specified in the prompt. However, we also noticed that the outputs contained many semantic errors (even after emphasizing the focus on semantic accuracy in the prompt, see Figure 2). This observation led us to evaluate the model outputs using word-level annotations focused on semantic accuracy errors (see §4).
@@ 3.4 Final Experiments
Taking the observations in §3.3 into account, we proceeded to generate the outputs on the test set of Quintd-1 for word-level error analysis. We first preprocessed the data as mentioned: we stripped out unnecessary fields, added units, and downsampled the data to fit the context. For all the models mentioned in §3.2, we used the prompt in Figure 2 and deterministic decoding with a maximum length of 512 tokens.
For comparison, we also generated outputs for the same inputs and identical prompts with GPT-3.5. Note that even though we fixed the temperature and seed to 00, the rest of the decoding parameters are inaccessible to us and may differ from the parameters we used for the open models.
@@ 4 Evaluation
For evaluation, we focus on identifying semantic errors in model outputs. We compare the generated texts to the input data, looking for parts of texts that are not faithful to the input data. We annotate the errors on the word level, considering all the words in the output text as potential sources of errors.
We use two complementary referenceless evaluation methods:
  • $\mathcal{E}_{\text{hum}}$: human evaluation based on crowdsourcing (§4.1),
  • $\mathcal{E}_{\text{gpt}}$: an automatic metric based on GPT-4 (§4.2).
The methods use similar instructions and produce outputs with equivalent format. The main idea is to compensate for the shortcomings of each approach: while human evaluation is costly (about ten times more expensive than automatic evaluation), using only an automatic metric based on a closed LLM would make the evaluation potentially non-reproducible and biased Kocmi and Federmann (2023a); Wang et al. (2023b). Reporting the results of both methods should hopefully increase the robustness and replicability of our results.
Our error taxonomy and its notation are inspired by Thomson and Reiter (2020) and Thomson et al. (2023). After preliminary examination of the outputs, we settled on four error categories: INCORRECTI, NOT_CHECKABLENC, MISLEADINGM, and OTHERO. To set clear boundaries between the categories, simplify the annotation interface and reach better inter-annotator agreement, we decided to keep our taxonomy more high-level than Thomson and Reiter and not to distinguish between fine-grained categories (e.g., incorrect name vs. incorrect number). The descriptions of our error categories, as presented in the instructions for annotation, are included in Table 2.
@@
INCORRECTI The fact in the text contradicts the data.
NOT_CHECKABLENC The fact in the text cannot be checked given the data.
MISLEADINGM The fact in the text is misleading in the given context.
OTHERO The text is problematic for another reason, e.g., grammatically or stylistically incorrect, irrelevant, or repetitive.
@@
data Nokia 3310 | color: black, blue, grey | display: 320x240px
text Nokia 3310 is produced in FinlandNC and features a 320x320I display. It is available in black colorM. The data seem to provide only partial information about the phone.O
Table 2: Categories of errors annotated in our evaluation and an example demonstrating the error types. See Appendix B for an explanation of individual errors in the example.
4.1 Human-based Evaluation
For the human annotation metric, we prepared a custom web interface, where an annotator is instructed to annotate text spans with respective error categories. We created custom visualizations for each data format (see Figure 3 and Appendix E for examples). [^8] We open-sourced our annotation framework as a stand-alone software package, see https://github.com/kasnerz/factgenie.
We hired annotators on the Prolific [^9] https://prolific.com crowdsourcing platform. In total, we hired 100 annotators, each annotating 20 examples (four model outputs for each of the five domains). We selected annotators with at least 10 completed tasks, a 100% approval rate and English as their primary language. We paid the annotators £9 per hour, according to the platform’s recommendations. The median time for completing the annotations was 47 minutes. See Appendix B for the instructions for the annotators and the annotation interface.